# {tocentry}

### v1

@@ 3.4 Chorismate Mutase
We equilibrated the structure from Ray et al. using our QM(S+R90)/MM MD for 12 ps at 300 K, with a 1 fs time step, at both the PBE/6-31G** and $\omega$B97X-3c levels of theory. The last frame of the PBE equilibration was used to initiate a 15-ps-long steered MD (SMD), biasing several inter-atomic distances to flip the Arg90 residue to form a hydrogen bond between the chorismate enolpyruvyl carboxylic oxygen and the arginine $\epsilon$-nitrogen. The SMD was performed at the PBE/6-31G** level, and the detailed protocol can be found in the SI. The resulting structure with a flipped Arg90 forms two hydrogen bonds (HBs) with the chorismate while the structure taken from Ref.61 features one chorismate-Arg90 hydrogen bond.
@@
For each combination of Arg90 conformation, QM region definition, and level of quantum theory, 11 configurations were selected from the last 10 ps of equilibration to initiate on-the-fly probability enhanced sampling (OPES) flooding simulations62. The configurations after the first 2-ps of each equilibration were used for initiating an SMD simulation to estimate the reaction free energy profile. The reaction coordinate (RC) $\xi$ for both OPES flooding and SMD was the bond-making C1--C9 distance minus the bond-breaking C3--ester oxygen distance. A harmonic restraint with a 1000 kJ/mol/Å2 force constant was used to drive the RC from its initial reactant value ($\xi\approx$1.8 Å) to -0.9 Å over 7.5 ps of SMD. The free energy surface obtained from the SMD was used to determine the barrier height parameter and the bias-excluded region defined as $\xi<\xi^{\mathrm{excl}}$ for the following OPES flooding. Detailed OPES flooding protocols are provided in the SI. All the enhanced sampling simulations (SMD and OPES flooding) used a 1 fs time step, with the enhanced sampling functionality provided by the PLUMED library63, 64.
@@
We also performed geometry optimization and nudged elastic band (NEB) simulations to find reaction transition states starting from both the 1HB and 2HB Arg90 conformations. The first frames of the $\omega$B97X-3c SMD with 1HB/2HB Arg90 and the S+R90 QM region were energy minimized at the same level of theory to produce local minimum energy reactant configurations. The last frames of the $\omega$B97X-3c SMD were geometrically aligned to the reactant structures, and their product-like active-site conformations were combined with the geometries in the reactant state for the remaining part of the system. The combined conformations were energy minimized to find the product state minima. Eight images were interpolated between the resulting reactant and product states using the image dependent pair potential method65. NEB optimization was performed without a climbing image, targeting a maximum force of 0.15 eV/Å, and then with a climbing image using the same convergence criterion. The geometry minimization and NEB optimization were performed on the QM/MM PES with the S+R90 QM region and the $\omega$B97X-3c functional, with the atoms of S+R90+R7+E78 allowed to move. All energy minimization was performed using i-PI and all NEB calculations were carried out using the Atomic Simulation Environment (ASE)66.
@@
We extracted the substrate and Arg90 geometries from the NEB conformations and performed single-point energy calculations in the gas phase. The energies were computed using PBE/6-31G** and $\omega$B97X-3c on every NEB image, and DLPNO-CCSD(T) at five geometries: the reactant, the product, the transition state and the two adjacent NEB images to the transition state. We performed the DLPNO-CCSD(T) calculations with cc-pVTZ and cc-pVQZ basis sets and used the two-point extrapolation formula67 to extrapolate both the HF and correlation energies to the complete basis set (CBS) limit. All DLPNO calculations were performed in ORCA68 with the “TightPNO” and “TightSCF” settings turned on. To reduce the error of the underlying quantum theory used in our MD simulations, we then refined the exchange parameters of the $\omega$B97X-3c functional to reproduce the DLPNO reaction energy barriers in both the 1HB and 2HB Arg90 conformations, optimizing both the fraction of short-range and of long-range exchange to exactly reproduce the two reaction barriers. We then checked to make sure that the resulting refined $\omega$B97X-3c functional also improved the energetics on the non-training points as well. We observed more accurate energies and curvature around the transition state, noting that these are related to the probability of successful barrier crossing in classical rate theory69.
@@ 4.1 Error Control via Charge-Octupole Estimation
Figure 2: QM/MM analysis for a QM-water-in-MM-water system. (A) The configuration of the system where QM/MM calculations were performed. The QM water molecules are shown as spheres and the MM water molecules are shown as sticks. (B) The error of QM/MM-Multipole (up-to-quadrupoles) QM-MM energy as a function of the cutoff radius $R_{\mathrm{cut}}$. (C) Mulliken charges of the QM atoms after up to 200 cycles of SCF with QM/MM-Ewald or QM/MM-Multipole using two $R_{\mathrm{cut}}$ values.
As discussed earlier, we truncate the multipole expansion at the quadrupole order (second order in the Taylor expansion), and thus the leading error in our approximation is the missing charge-octupole (the third order) interaction. This suggests that the interaction energy between the region II charges and the QM octupoles can serve as an error estimate,
@@
This estimate assumes that region I is sufficiently large that any higher-order terms (fourth order and above) are negligible. From here on, we consider region I to be defined by a sphere of radius $R_{\mathrm{cut}}$ centered at the QM subsystem geometric center.
@@
To validate the charge-octupole (octupole for short) error estimate, we examined a simple QM-water-in-MM-water system where computing the exact QM-MM electrostatic coupling energy is feasible. The system consists of 7 QM water molecules embedded in 221 MM water molecules (Figure 2A). The exact QM-MM electrostatic coupling was computed by the standard Ewald method, treating the electron density on the DFT grid as one group of point charges and the MM water molecules as another. The electron density used in the test was obtained from a converged QM/MM-Multipole SCF with $R_{\mathrm{cut}}=$15.5 Å. In Figure 2B, we show the exact error of including up to quadrupoles ($E_{\mathrm{exact}}-E_{\mathrm{approx}}$) versus the region II octupole error estimate as a function of $R_{\mathrm{cut}}$. (We did not re-optimize the density with different $R_{\mathrm{cut}}$ in this test, thus the $R_{\mathrm{cut}}$ dependency is entirely due to the varying number of exactly computed QM-MM interactions). We find a good agreement between the octupole estimate and the exact error beyond $R_{\mathrm{cut}}\approx 12.5$. For a smaller $R_{\mathrm{cut}}$, the octupole error estimate remains qualitatively correct, and is of the same order of magnitude as the exact error.
@@
Figure 3: QM/MM analysis of a model for microtubule-mediated GTP hydrolysis. (A) The system configuration used for the QM/MM calculations. The QM atoms are shown as spheres, the protein is depicted as ribbons (with subunit $\alpha$ in blue and subunit $\beta$ in red), and the rest of the system is represented by transparent surfaces. (B) A zoomed-in view of the QM region. The QM atoms are shown as balls and sticks, the MM part of GTP is represented by thin sticks, and the protein is shown in transparent blue ($\alpha$ subunit) and red ($\beta$ subunit). The atoms are colored as follows: hydrogen (white), oxygen (red), nitrogen (blue), carbon (gray), magnesium (pink), and phosphorus (gold). (C) The error in the QM/MM-Multipole SCF-converged energy as a function of the cutoff radius $R_{\text{cut}}$, assuming the energy obtained with $R_{\text{cut}}=40$ is the ground truth. The octupole error estimate was computed using the density converged with a cutoff of $R^{\prime}_{\text{cut}}$.
@@
We next examine a more complex system, a model of GTP hydrolysis in microtubules. The system composition and QM subsystem definition are detailed in the Computational Details section as well as shown in Figure 3A and Figure 3B. The electronic structure was described at the $\omega$B97X-3c level in this analysis. We estimated the exact error by comparing the energy at $R_{\mathrm{cut}}=40$ (which we take as exact) with the SCF-converged energy $E(R_{\mathrm{cut}})$ at various $R_{\mathrm{cut}}$ values. Unlike in the previous example, this error depends on $R_{\mathrm{cut}}$ not only through the omission of octupole and higher interactions with region II but also through the $R_{\mathrm{cut}}$-dependency in the converged QM density, from the absence of polarization effects at the above-quadrupole level by region II charges. According to the linear response theory of HF/KS70, the error in the density is proportional to the error in the electronic Hamiltonian. As such, the lack of octupole polarization by a charge at distance $R$ introduces a density error that scales in the same way as the charge-octupole interaction ($\propto 1/R^{4}$). Integrating over the whole region II, this gives an accumulated error scaling as $\int_{R_{\mathrm{cut}}}^{\infty}4\pi R^{2}/R^{4}\mathrm{d}R\propto 1/R_{\mathrm{cut}}$. Due to the variational condition, the total energy error scales quadratically with the density error, yielding a $1/R_{\mathrm{cut}}^{2}$ decay, faster than the $1/R_{\mathrm{cut}}$ scaling of the total octupole interaction with region II. Hence, the leading term of the error at large $R_{\mathrm{cut}}$ is still the missing octupole interaction, the same as in the previous test with a frozen density. As shown in Figure 3C, the octupole error estimate obtained by simply summing the missing charge-octupole interactions agrees well with the “exact” error at sufficiently large $R_{\mathrm{cut}}$. Furthermore, since the energy error due to the density error decays quickly ($1/R_{\mathrm{cut}}^{2}$), the octupole error estimate may be computed using a density converged with a small region I (defined by $R_{\mathrm{cut}}^{\prime}$), and we show this is indeed the case in Figure 3.
@@ 4.2 Timings
For the PBE calculations, the pseudo-bond parameters for $\omega$B97X-V were used to avoid building another parametrization. The wall time per MD step was averaged over 100 MD steps, with a fixed $R_{\mathrm{cut}}=25$Å. The MT system contains 156 QM atoms, and the CM system has 43 QM atoms in the S+R90 definition of the QM subsystem and 72 QM atoms in the S+R90+R7+E78 definition of the QM subsystem. The number of computational basis functions ($N_{\mathrm{AO}}$) is 1756, 532, and 846 using the 6-311G** basis in the MT and for the two CM QM subsystem definitions, respectively. The $\omega$B97X-3c functional comes with a specially parameterized double-zeta basis, and $N_{\mathrm{AO}}=1309$, 398 and 637 in the MT and two CM QM subsystem definitions. The timing results are provided in Table 1.
System PBE/6-311G** $\omega$B97X-3c $\omega$B97X-V/6-311G**
MT 1900$\pm$90 730$\pm$20 370$\pm$20
CM (S+R90) 11200$\pm$600 6600$\pm$300 4400$\pm$200
CM (S+R90+R7+E78) 6900$\pm$300 3290$\pm$80 1790$\pm$80
Table 1: Computational performance of the QM/MM MD simulations using 32 CPU cores (AMD EPYC 7742) and one A100 GPU. The reported values are the mean number of steps per day, with the fluctuations computed as the standard deviation over 100 MD steps.
We observe that in our current implementation of periodic ab initio QM/MM MD, simulations with a hybrid density functional ($\omega$B97X-3c) on systems with $>100$ atoms ($>1000$ atomic basis functions) can be run with a throughput of $\sim 1000$ MD steps per day on a single GPU. Although a careful comparison will require controlling many other details and will be considered in a later study, this throughput appears to be roughly comparable (e.g. within a factor of $\sim 2$) to that reported by recent GPU implementations of non-periodic ab initio QM/MM MD71.
@@ 4.3 SCF Stability
If one ignores the dipole and quadrupole terms, the QM/MM-Multipole approach reduces to the QM/MM-Ewald approach using Mulliken charges20. It is known that the QM/MM-Ewald method with Mulliken charges suffers from instabilities in the SCF optimization when diffuse basis functions are used. The instability is related to the definition of the Mulliken charges, which becomes increasingly unsatisfactory in diffuse bases (although a description of the electrostatic potential at long distances is still provided by the full set of Mulliken multipoles). In the test system shown in Figure 2A, we show that we could not converge the SCF within 200 cycles using the QM/MM-Ewald approach even when $R_{\mathrm{cut}}$ was set to half the box length. After extending region I beyond the unit cell, with $R_{\mathrm{cut}}=$14 Å, we were able to finally converge the QM/MM-Ewald SCF. The converged electron density, however, still has large errors revealed by large positive/negative Mulliken charges on the oxygen/hydrogen atoms. In contrast, the QM/MM-Multipole approach is always stable and produced reasonable Mulliken charges, even with the smaller $R_{\mathrm{cut}}$.

### v2

@@ 3.4 Chorismate Mutase
We equilibrated the structure from Ray et al. using our QM(S+R90)/MM MD for 12 ps at 300 K, with a 1 fs time step, at both the PBE/6-31G** and $\omega$B97X-3c levels of theory. The last frame of the PBE equilibration was used to initiate a 15-ps-long steered MD (SMD), biasing several inter-atomic distances to flip the Arg90 residue to form a hydrogen bond between the chorismate enolpyruvyl carboxylic oxygen and the arginine $\epsilon$-nitrogen. The SMD was performed at the PBE/6-31G** level, and the detailed protocol can be found in the SI. The resulting structure with a flipped Arg90 forms two hydrogen bonds (HBs) with the chorismate while the structure taken from Ref.79 features one chorismate-Arg90 hydrogen bond.
@@
For each combination of Arg90 conformation, QM region definition, and level of quantum theory, 11 configurations were selected from the last 10 ps of equilibration to initiate on-the-fly probability enhanced sampling (OPES) flooding simulations80. The configurations after the first 2-ps of each equilibration were used for initiating an SMD simulation to estimate the reaction free energy profile. The reaction coordinate (RC) $\xi$ for both OPES flooding and SMD was the bond-making C1--C9 distance minus the bond-breaking C3--ester oxygen distance. A harmonic restraint with a 1000 kJ/mol/Å2 force constant was used to drive the RC from its initial reactant value ($\xi\approx$1.8 Å) to -0.9 Å over 7.5 ps of SMD. The free energy surface obtained from the SMD was used to determine the barrier height parameter and the bias-excluded region defined as $\xi<\xi^{\mathrm{excl}}$ for the following OPES flooding. Detailed OPES flooding protocols are provided in the SI. All the enhanced sampling simulations (SMD and OPES flooding) used a 1 fs time step, with the enhanced sampling functionality provided by the PLUMED library81, 82.
@@
We also performed geometry optimization and nudged elastic band (NEB) simulations to find reaction transition states starting from both the 1HB and 2HB Arg90 conformations. The first frames of the $\omega$B97X-3c SMD with 1HB/2HB Arg90 and the S+R90 QM region were energy minimized at the same level of theory to produce local minimum energy reactant configurations. The last frames of the $\omega$B97X-3c SMD were geometrically aligned to the reactant structures, and their product-like active-site conformations were combined with the geometries in the reactant state for the remaining part of the system. The combined conformations were energy minimized to find the product state minima. Eight images were interpolated between the resulting reactant and product states using the image dependent pair potential method83. NEB optimization was performed without a climbing image, targeting a maximum force of 0.15 eV/Å, and then with a climbing image using the same convergence criterion. The geometry minimization and NEB optimization were performed on the QM/MM PES with the S+R90 QM region and the $\omega$B97X-3c functional, with the atoms of S+R90+R7+E78 allowed to move. All energy minimization was performed using i-PI and all NEB calculations were carried out using the Atomic Simulation Environment (ASE)84.
@@
We extracted the substrate and Arg90 geometries from the NEB conformations and performed single-point energy calculations in the gas phase. The Arg90 sidechain was capped by a hydrogen positioned along the Cα-Cβ direction at a distance of 1.08 Å from Cβ. The energies were computed using PBE/6-31G** and $\omega$B97X-3c on every NEB image, and LNO-CCSD(T)85 at five geometries: the reactant, the product, the transition state, and the two adjacent NEB images to the transition state. We performed the LNO-CCSD(T) calculations with cc-pVTZ and cc-pVQZ basis sets and used the two-point extrapolation formula86 to extrapolate both the HF and correlation energies to the complete basis set (CBS) limit. All LNO calculations were performed with our recent implementation87 in PySCFAD88. A natural occupation threshold of $10^{-5}$ was chosen for truncating the local virtual natural orbitals, while a natural occupation threshold of $10^{-4}$ was used for truncating the occupied natural orbitals. Such a choice of thresholds was found to give 0.05 kcal/mol mean absolute error in the LNO-CCSD(T)/cc-pVDZ energy on the five 1HB NEB geometries and 0.02 kcal/mol on the 2HB NEB geometries, compared to the canonical CCSD(T)/cc-pVDZ results. All the cc-pVDZ, cc-pVTZ, and cc-pVQZ calculations were also found to be converged to within 0.06 kcal/mol in MAE with the $10^{-4}$ occupied/$10^{-5}$ virtual occupancy thresholds when comparing to the energies with larger thresholds ($2\times 10^{-4}$ occupied/$2\times 10^{-5}$ virtual). To reduce the error of the underlying quantum theory used in our MD simulations, we then refined the exchange parameters of the $\omega$B97X-3c functional to reproduce the LNO-CCSD(T)/CBS reaction energy barriers in both the 1HB and 2HB Arg90 conformations, optimizing both the fraction of short-range and of long-range exchange to exactly reproduce the two reaction barriers. We then checked to make sure that the resulting refined $\omega$B97X-3c functional also improved the energetics on the non-training points as well. We observed more accurate energies and curvature around the transition state, noting that these are related to the probability of successful barrier crossing in classical rate theory89.
@@ 4.1 Error Control via Charge-Octupole Estimation
Figure 2: QM/MM analysis for a QM-water-in-MM-water system. (A) The configuration of the system where QM/MM calculations were performed. The QM water molecules are shown as spheres and the MM water molecules are shown as sticks. (B) The error of QM/MM-Multipole (up-to-quadrupoles) QM-MM energy as a function of the region I cutoff radius $R_{\mathrm{cut}}$. (C) Mulliken charges of the QM atoms after up to 200 cycles of SCF with QM/MM-Ewald or QM/MM-Multipole using two $R_{\mathrm{cut}}$ values.
As discussed earlier, we truncate the multipole expansion at the quadrupole order (second order in the Taylor expansion) for the electrostatics between region II and the QM density in the reference cell, and thus the leading error in our approximation is the missing charge-octupole (the third order) interaction. This suggests that the interaction energy between the region II charges and the QM octupoles can serve as an error estimate,
@@
This estimate assumes that region I is sufficiently large that any higher-order terms (fourth order and above) are negligible. From here on, we consider region I to be defined by a sphere of radius $R_{\mathrm{cut}}$ centered at the QM subsystem geometric center. This cutoff truncates the short-range interactions of the QM octupoles and higher order multipoles with the charges outside region I, and should not be confused with the Ewald real-space cutoff.
@@
To validate the charge-octupole (octupole for short) error estimate, we examined a simple QM-water-in-MM-water system where computing the exact QM-MM electrostatic coupling energy is feasible. The system consists of 7 QM water molecules embedded in 221 MM water molecules (Figure 2A). The exact QM-MM electrostatic coupling was computed by the standard Ewald method, treating the electron density on the DFT grid as one group of point charges and the MM water molecules as another. The electron density used in the test was obtained from a converged QM/MM-Multipole SCF with $R_{\mathrm{cut}}=$15.5 Å. In Figure 2B, we show the exact error of including up to quadrupoles ($E_{\mathrm{exact}}-E_{\mathrm{approx}}$) versus the region II octupole error estimate as a function of $R_{\mathrm{cut}}$. (We did not re-optimize the density with different $R_{\mathrm{cut}}$ in this test, thus the $R_{\mathrm{cut}}$ dependency is entirely due to the varying number of exactly computed QM-MM interactions). We find a good agreement between the octupole estimate and the exact error beyond $R_{\mathrm{cut}}\approx 12.5$. For a smaller $R_{\mathrm{cut}}$, the octupole error estimate remains qualitatively correct and is of the same order of magnitude as the exact error.
@@
Figure 3: QM/MM analysis of a model for microtubule-mediated GTP hydrolysis. (A) The system configuration used for the QM/MM calculations. The QM atoms are shown as spheres, the protein is depicted as ribbons (with subunit $\alpha$ in blue and subunit $\beta$ in red), and the rest of the system is represented by transparent surfaces. (B) A zoomed-in view of the QM region. The QM atoms are shown as balls and sticks, the MM part of GTP is represented by thin sticks, and the protein is shown in transparent blue ($\alpha$ subunit) and red ($\beta$ subunit). The atoms are colored as follows: hydrogen (white), oxygen (red), nitrogen (blue), carbon (gray), magnesium (pink), and phosphorus (gold). (C) The error in the QM/MM-Multipole SCF-converged energy as a function of the region I cutoff radius $R_{\text{cut}}$, assuming the energy obtained with $R_{\text{cut}}=40$ is the ground truth. The octupole error estimate was computed using the density converged with a cutoff of $R^{\prime}_{\text{cut}}$.
@@
We next examine a more complex system, a model of GTP hydrolysis in microtubules. The system composition and QM subsystem definition are detailed in the Computational Details section as well as shown in Figure 3A and Figure 3B. The electronic structure was described at the $\omega$B97X-3c level in this analysis. We estimated the exact error by comparing the energy at $R_{\mathrm{cut}}=40$ (which we take as exact) with the SCF-converged energy $E(R_{\mathrm{cut}})$ at various $R_{\mathrm{cut}}$ values. Unlike in the previous example, this error depends on $R_{\mathrm{cut}}$ not only through the omission of octupole and higher interactions with region II but also through the $R_{\mathrm{cut}}$-dependency in the converged QM density, from the absence of polarization effects at the above-quadrupole level by region II charges. According to the linear response theory of HF/KS90, the error in the density is proportional to the error in the electronic Hamiltonian. As such, the lack of octupole polarization by a charge at distance $R$ introduces a density error that scales in the same way as the charge-octupole interaction ($\propto 1/R^{4}$). Integrating over the whole region II, this gives an accumulated error scaling as $\int_{R_{\mathrm{cut}}}^{\infty}4\pi R^{2}/R^{4}\mathrm{d}R\propto 1/R_{\mathrm{cut}}$. Due to the variational condition, the total energy error scales quadratically with the density error, yielding a $1/R_{\mathrm{cut}}^{2}$ decay, faster than the $1/R_{\mathrm{cut}}$ scaling of the total octupole interaction with region II. Hence, the leading term of the error at large $R_{\mathrm{cut}}$ is still the missing octupole interaction, the same as in the previous test with a frozen density. As shown in Figure 3C, the octupole error estimate obtained by simply summing the missing charge-octupole interactions agrees well with the “exact” error at sufficiently large $R_{\mathrm{cut}}$. Furthermore, since the energy error due to the density error decays quickly ($1/R_{\mathrm{cut}}^{2}$), the octupole error estimate may be computed using a density converged with a small region I (defined by $R_{\mathrm{cut}}^{\prime}$), and we show this is indeed the case in Figure 3.
@@ 4.2 Timings
For the PBE and B3LYP calculations, the pseudo-bond parameters for $\omega$B97X-V were used to avoid building another parametrization. The wall time per MD step was averaged over 100 MD steps, with a fixed $R_{\mathrm{cut}}=25$Å. The MT system contains 156 QM atoms, and the CM system has 43 QM atoms in the S+R90 definition of the QM subsystem and 72 QM atoms in the S+R90+R7+E78 definition of the QM subsystem. The number of computational basis functions ($N_{\mathrm{AO}}$) is 1756, 532, and 846 using the 6-311G** basis in the MT and for the two CM QM subsystem definitions, respectively, and $N_{\mathrm{AO}}$=1513, 475, and 744 for the 6-311G* basis set. The $\omega$B97X-3c functional comes with a specially parameterized double-zeta basis, and $N_{\mathrm{AO}}=1309$, 398 and 637 in the MT and two CM QM subsystem definitions. The timing results are provided in Table 1.
System PBE/6-311G** B3LYP/6-311G* $\omega$B97X-3c $\omega$B97X-V/6-311G**
MT 1900$\pm$90 1510$\pm$70 730$\pm$20 370$\pm$20
CM (S+R90) 11200$\pm$600 9900$\pm$500 6600$\pm$300 4400$\pm$200
CM (S+R90+R7+E78) 6900$\pm$300 5800$\pm$300 3290$\pm$80 1790$\pm$80
Table 1: Computational performance of the QM/MM MD simulations using 32 CPU cores (AMD EPYC 7742) and one A100 GPU. The reported values are the mean number of steps per day, with the fluctuations computed from the standard deviation over 100 MD steps.
We observe that in our current implementation of periodic ab initio QM/MM MD, simulations with a hybrid density functional ($\omega$B97X-3c) on systems with $>100$ atoms ($>1000$ atomic basis functions) can be run with a throughput of $\sim$1000 MD steps per day on a single GPU. Although a careful comparison will require controlling many other details and will be considered in a later study, the speed of B3LYP/6-311G* on the MT system appears to be roughly comparable to that reported for B3LYP on a system with a similar size in recent GPU implementations of non-periodic ab initio QM/MM MD91.
@@ 4.3 SCF Stability
If one ignores the dipole and quadrupole terms, the QM/MM-Multipole approach reduces to the QM/MM-Ewald approach using Mulliken charges23. (More precisely, the original QM/MM-Ewald approach additionally fixes region I to be the unit cell but we allow a flexible choice of region I herein). It is known that the QM/MM-Ewald method with Mulliken charges suffers from instabilities in the SCF optimization when diffuse basis functions are used. The instability is related to the definition of the Mulliken charges, which becomes increasingly unsatisfactory in diffuse bases (although a description of the electrostatic potential at long distances is still provided by the full set of Mulliken multipoles). In the test system shown in Figure 2A, we show that we could not converge the SCF within 200 cycles using the QM/MM-Ewald approach even when $R_{\mathrm{cut}}$ was set to half the box length. After extending region I beyond the unit cell, with $R_{\mathrm{cut}}=$14 Å, we were able to finally converge the QM/MM-Ewald SCF. The converged electron density, however, still has large errors revealed by large positive/negative Mulliken charges on the oxygen/hydrogen atoms. In contrast, the QM/MM-Multipole approach is always stable and produces reasonable Mulliken charges, even with the smaller $R_{\mathrm{cut}}$.